OpenStudy (ellie202000):

For each function shown: 1) Name the starting point. 2) Explain how y changes as x increases. Function #1: y = 0.5 + 3x Function #2: y = 3(0.5)x

OpenStudy (ellie202000):

@welshfella

OpenStudy (welshfella):

1) starting point is where x = 0 y = 0.5 + 3(0) as x increases the value of y will also increase

OpenStudy (welshfella):

2) starting point= 3(0.5)* 0 = ? what do you think happens to y as x increases?

OpenStudy (ellie202000):

As x increase y would too?

OpenStudy (ellie202000):

Would the starting point for function #1 be .5 and the starting point for function #2 be 1?